# Training Large Neural Networks With Low-Dimensional Error Feedback

### v1

@@ Appendix B Learning rule implementation
Our implementation and optimization were conducted entirely using PyTorch. For the RAF layers, we initialized all weights using Kaiming uniform initialization. We modified the backward pass by adjusting the gradients with respect to the input, ensuring they align with our proposed update rule. In the output layer, we learn the projection matrix $\boldsymbol{P}$ to capture the principal directions of the target labels $\boldsymbol{y}$. In the hidden layers, $\boldsymbol{P}$ is learned to project onto the principal directions of the error signal from the subsequent layer, represented as $\boldsymbol{\delta}_{l+1}$. Specifically, for a layer $l$ with input dimension $n$, output dimension $m$, and rank constraint $r$, we proceed as follows:
@@
Algorithm 1 Modified Backward Pass for RAF Layer
The derivation of the dRAF (directed restricted adaptive feedback) follows the same procedure, with the key difference being that the error signal $\boldsymbol{\delta}$ does not necessarily originate from the next layer; instead, it can come from any subsequent layer.
@@ Appendix C Numerical experiments
In Figure [3.a], we present the results of training a network with four hidden layers, each containing 512 neurons, and an output layer with 10 neurons on the CIFAR-10 dataset. The network was trained using RAF, as described earlier, without rank constraints, except for one layer at a time. For comparison, we also trained the network using standard backpropagation as a baseline. All networks were trained with a batch size of 32, a learning rate of $6\times 10^{-4}$, and weight decay of $4\times 10^{-4}$. The Adam optimizer with AMSGrad was used, with training conducted for 160 epochs and an exponential learning rate decay factor of 0.975. Each experiment was repeated 10 times.
@@
dRAF
For the results shown in Figure [3.d], we trained the same model using dRAF. In one experiment, we propagated the error signal from the last layer to all preceding layers. In a separate experiment, we propagated the error signal directly from the penultimate layer to all earlier layers, applying rank constraints to these layers while keeping the last layer at full rank. All optimizations were performed with a batch size of 32, a learning rate of $6\times 10^{-4}$, and weight decay of $4\times 10^{-4}$. The Adam optimizer with AMSGrad was used, training for 160 epochs with an exponential learning rate decay factor of 0.975. Each training run was repeated 5 times
@@
To extend our RAF algorithm to convolutional layers, we apply the rank constraint to the number of channels in the error signal. This effectively constrains the dimensionality of the error signal across the spatial dimensions. Similar to the implementation for fully connected layers in Algorithm 1, we modify the backward pass in the same way for convolutional layers.
@@
Training was performed using the Restricted Adaptive Feedback (RAF) method, where rank constraints were applied to feedback sent to the retina with ranks $r=2,4,32$. The network was trained with hyperparameters similar to those in Linsdey et. al. [29], using the RMSProp optimizer with a learning rate of $1\times 10^{-4}$, a weight decay of $1\times 10^{-5}$, and an exponential learning rate decay factor of 0.985. Each model was trained for 120 epochs, and all experiments were repeated five times to ensure robustness.
Additionally, we conducted experiments with dynamic Restricted Adaptive Feedback (dRAF), where feedback to the retina originated from higher visual layers while feedforward connections within the retina were constrained to four channels.

### v2

@@ Appendix B Learning rule implementation
Our implementation and optimization were conducted entirely using PyTorch. For the LDFA layers, we initialized all weights using Kaiming uniform initialization. We modified the backward pass by adjusting the gradients with respect to the input, ensuring they align with our proposed update rule. In the output layer, we learn the projection matrix $\boldsymbol{P}$ to capture the principal directions of the target labels $\boldsymbol{y}$. In the hidden layers, $\boldsymbol{P}$ is learned to project onto the principal directions of the error signal from the subsequent layer, represented as $\boldsymbol{\delta}_{l+1}$. Specifically, for a layer $l$ with input dimension $n$, output dimension $m$, and rank constraint $r$, we proceed as follows:
@@
Algorithm 1 Modified Backward Pass for LDFA Layer
The derivation of the dLDFA (directed restricted adaptive feedback) follows the same procedure, with the key difference being that the error signal $\boldsymbol{\delta}$ does not necessarily originate from the next layer; instead, it can come from any subsequent layer.
@@ Appendix C Numerical experiments
In Figure [3.a], we present the results of training a network with four hidden layers, each containing 512 neurons, and an output layer with 10 neurons on the CIFAR-10 dataset. The network was trained using LDFA, as described earlier, without rank constraints, except for one layer at a time. For comparison, we also trained the network using standard backpropagation as a baseline. All networks were trained with a batch size of 32, a learning rate of $6\times 10^{-4}$, and weight decay of $4\times 10^{-4}$. The Adam optimizer with AMSGrad was used, with training conducted for 160 epochs and an exponential learning rate decay factor of 0.975. Each experiment was repeated 10 times.
@@
Direct LDFA
For the results shown in Figure [3.d], we trained the same model using dLDFA. In one experiment, we propagated the error signal from the last layer to all preceding layers. In a separate experiment, we propagated the error signal directly from the penultimate layer to all earlier layers, applying rank constraints to these layers while keeping the last layer at full rank. All optimizations were performed with a batch size of 32, a learning rate of $6\times 10^{-4}$, and weight decay of $4\times 10^{-4}$. The Adam optimizer with AMSGrad was used, training for 160 epochs with an exponential learning rate decay factor of 0.975. Each training run was repeated 5 times
@@
To extend our LDFA algorithm to convolutional layers, we apply the rank constraint to the number of channels in the error signal. This effectively constrains the dimensionality of the error signal across the spatial dimensions. Similar to the implementation for fully connected layers in Algorithm 1, we modify the backward pass in the same way for convolutional layers.
@@
Training was performed using the low-dimensional feedback alignment (LDFA) method, where rank constraints were applied to feedback sent to the retina with ranks $r=2,4,32$. The network was trained with hyperparameters similar to those in Linsdey et. al. [29], using the RMSProp optimizer with a learning rate of $1\times 10^{-4}$, a weight decay of $1\times 10^{-5}$, and an exponential learning rate decay factor of 0.985. Each model was trained for 120 epochs, and all experiments were repeated five times to ensure robustness.
Additionally, we conducted experiments with dynamic Restricted Adaptive Feedback (dLDFA), where feedback to the retina originated from higher visual layers while feedforward connections within the retina were constrained to four channels.